• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name epstein, templon

    Log entry time 11:48:55 on December 18,1999

    Entry number 31400

    keyword=config change, kin 27 production

    Finished the cross normalization run with Spec_e at 87.388 degrees and hadron arm left fixed at the elastic scattering and kin. 17 settings. Now changing hadron arm to kin. #27 settings. Starting the hadron Q2 and Q3 cycling. Bringing current up to 1400 A in each. then waited five minutes, dropped by hand to 1000 A, finally
    let control system have it and bring to desired p0 value.

    also changed hadron angle to 35.38 deg. target was 35.39 deg but our
    vernier landed at -54 instead of -53.3, blame Templon for this error.

    right now we are waiting for the hadron dipole to drift on down.